OpenStudy (anonymous):

Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3. −21 −16 −10 −1

OpenStudy (bibby):

replace f(x) with y, and exchange variables 4(x+3) − 5 = y 4(y+3) - 5 = x

OpenStudy (bibby):

solve for y

OpenStudy (anonymous):

which choice? sorry to bother

OpenStudy (bibby):

dunno, do the math. Solve for y \[\huge 4(y+3) - 5 = x\]

OpenStudy (anonymous):

- 21 -16 -10 or -1

OpenStudy (anonymous):

ill give you a medal right now

OpenStudy (bibby):

I don't want a medal. I got more medals than I can ever use

OpenStudy (anonymous):

can u atleast help me for my good? which choice lol

OpenStudy (bibby):

"your good" = learning the math Add 5 to both sides, divide, subtract. plug in.

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

hmm got 17/4

OpenStudy (bibby):

I got a different answer

OpenStudy (anonymous):

u sure its right lol

OpenStudy (anonymous):

add 5 to 12 is what i did

OpenStudy (bibby):

I added 5 to both sides of \[\huge 4(y+3) - 5 = x\]and got\[\huge 4(y+3) = x + 5\]

OpenStudy (the_fizicx99):

Bib you're too good :"D

OpenStudy (zzr0ck3r):

the answer would be the solution to f(x) = 3